OpenStudy (anonymous):

How would you solve this without a calculator? cos (tan^-1(2/3))

OpenStudy (alekos):

start by drawing a right triangle with one angle theta, adjacent side 3 and opposite 2 i.e. tan(theta) = 2/3 got that?

OpenStudy (anonymous):

Yes, then you do the pythagorean theorem to get the hypotenuse which is 13 right?

OpenStudy (alekos):

yes but its sqrt13. do you know what to do next?

OpenStudy (anonymous):

Ok. No what do I do next

OpenStudy (alekos):

ok, theta = tan^-1(2/3) so cos(tan^-1(2/3) = cos(theta) which you can work out from your triangle!

OpenStudy (anonymous):

I don't really understand

OpenStudy (anonymous):

Do you mean figure out tan^-1(2/3) first then whatever that answer is take the cos of?

OpenStudy (alekos):

if tan(theta) = 2/3 then theta = tan^-1(2/3) happy with that?

OpenStudy (anonymous):

So regardless you have to already know what tan^-1(2/3) is? Like have the unit circle memorized with the tangents?

OpenStudy (alekos):

no you dont have to memorise anything because we are not going to work out what theta is. do you follow what i did before?

OpenStudy (anonymous):



OpenStudy (alekos):

great diagram and theta will be on the bottom left. do you understand about theta = tan^-1(2/3)?

OpenStudy (anonymous):

The only thing I get from that is that you have to know that value

OpenStudy (anonymous):

Which I can figure out if I know the unit circle

OpenStudy (anonymous):

What do you mean by that?

OpenStudy (alekos):

no, you are stuck on thinking that you need to know the angle theta and you dont can you see that theta = tan^-1(2/3), because we have not finished yet

OpenStudy (anonymous):

Ok yes, so theta is still undefined

OpenStudy (alekos):

yes and it will stay that way now going back to the beginning cos(tan^-1(2/3)) = cos(theta) and we can work out cos(theta) from your triangle!

OpenStudy (anonymous):

Ok.. so cos is adjacent over hypotenuse, and that would be 3/sqrt13? Is that what you mean?